OpenStudy (anonymous):

(x-6)/(x^2 +4x-32) • (x-4)/(x+2) please help

OpenStudy (anonymous):

\[ \frac{(x-6) (x-4)}{(x+2) \left(x^2+4 x-32\right)}=\frac{(x-6) (x-4)}{(x+2) (x-4) (x+8)}=\frac{x-6}{(x+2) (x+8)} \]